kg and 2.159 x 26 are singly ionized (+e) and Two isotopes of carbon, carbon-12 and carbon-13, have masses of 1.993 x 10 kg, respectively. These two is each is given a speed of 7.38 x 105 m/s. The ions then enter the bending region of a mass spectrometer where the magnetic field is 0.5100 T. Determine the spatial separation between the two isotopes after they have traveled through a half-circle.
Answer #1

The spatial sepration of the ions is

d = \left ( \frac{2v}{Bq} \right )\left ( m_2-m_1 \right )

d = \left ( \frac{2(7.38 \times 10^5)}{(0.51)(1.6 \times 10^{19})} \right )\left (2.159 \times 10^{-26}-1.993 \times 10^{-26} \right )

d = 0.03 m